OpenStudy (moongazer):

The average of X numbers is 5. If Y is added to each number, the average will be? please explain (medal for a good answer)

OpenStudy (anonymous):

If the average of x numbers is 5 then the sum of those numbers is 5x. Example 2 3 4 5 6 7 8 have an average of 5 and a sum of 35 which is 7(5) so the sum would be 5x + y Then to find the average = you have x numbers that had an average of 5 plus you added one more number (y) to the sum so you now have x + 1 numbers Average is 5x + y ------ x + 1

OpenStudy (anonymous):

youre adding x*y though if you add y to x numbers so the total is 5x+xy. And the average is still of x numbers. So (5x+xy)/(x)=5+y

OpenStudy (anonymous):

You are correct... I read the problem wrong.. good one..